OpenStudy (anonymous):

Which expressions are equivalent to 4(4x + 2y + x − y)? 4(5x + 3y) and 20x + 12y 4(5x + y) and 20x + 4y 4(3x + y) and 12x + 4y 4(3x + 3y) and 12x + 12y

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

yea I am pretty bad myself

OpenStudy (anonymous):

what grade work is this

OpenStudy (anonymous):

@pooja195 @math1234

OpenStudy (anonymous):

6 (advanced) I do mostly 8th or 9th tho

OpenStudy (anonymous):

@horsegirl27 @HelpOfTheGods @Wolverine32 @omgitsjc

OpenStudy (anonymous):

The first expression on the left just requires you to combine the like terms on the inside of the parentheses.

OpenStudy (anonymous):

That would be\[4x + 2y +x - y = (4+1)x + (2-1)y = 5x + y\]

OpenStudy (anonymous):

so th answer is B

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Yay! can you help with another?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

What is the value of fraction 1 over 3x3 + 5.2y when x = 3 and y = 2? 13.4 16.4 19.4 28.4

OpenStudy (anonymous):